undeadknight26 (undeadknight26):

Is this right? 3. Two boys want to use a seesaw and they need to move the seesaw so that their weights will balance out. The formula is given by w1 • d1 = w2 • d2 where w1 = weight of the first boy, d1 = distance of the first boy from the fulcrum, w2 = weight of the second boy, and d2 = distance of the second boy from the fulcrum. Rewrite the formula to solve for d2. Show all steps in your work.

undeadknight26 (undeadknight26):

I put w^1 * d^1=w^2 * d^2 /w^2 /w^2 d2 = w1*d1/w2

OpenStudy (perl):



undeadknight26 (undeadknight26):

i think i did something wrong on the equation and left out the w's? thats what my teacher said.